Tech Meeting Minutes 20100520
Present: Graeme, Mike, Sam, Andrew, Wahid, Stuart
|Table of contents|
- LHCb upload problems at Glasgow. Thought this was completely solved, but there is still a persistent level of failures. Being tracked within UK. There is a definite issue with globus gridftp and NAT.
- LHCb problems at Edinburgh. Ticket 57042 was updated with the latest news. Asked LHCb to close and open new tickets for any outstanding issues.
- squid: ECDF have access to Glasgow and Lancaster squids. ACTION: Wahid to ask for update of firewall at ECDF to get to Lancaster. ACTION: Follow-up with Toa change request.
- CREAM: Have resources freed up at ECDF for this, hopefully some progress next week.
- Storm + GPFS: Problem found where ACL changes seem to be slow (few seconds) on nfs mounted gpfs (affects 1/20 transfers). Medium term issue with whether grid use of gpfs would be covered in ECDF's payments to IBM, but still hope to move towards this solution.
- Durham: Need to check site is fully upgraded.
- ATLAS pilot code which supports glexec is now checked in. ACTION: Graeme to ask Jose to check this now.
- Glasgow has had first evaluation meetings. Now going back for clarification on some points.
- Meeting next week: http://indico.cern.ch/conferenceDisplay.py?confId=95724.